# Quantum teleportation implies symmetry-protected topological order

### v2

@@ Appendix I Proof of Lemma 11
$=\bar{A}_{1,n,\nu}\cdots\bar{A}_{m,n,\nu}\,\sigma^{\nu}_{f_{n}}\,,~{}~{}$ (K.157)
where the first label on $\bar{A}_{s,n,\nu}$ corresponds to the “measurement region” $s\in[1,m]$ Friedman et al. (2022a), $n$ labels the logical qubit, and $\nu$ labels the logical basis operator (e.g., $\nu\in\{x,z\}$).
We now work out the optimal spacing between measurement sites $j_{s,n,\nu}$ for different $s$. Following the proofs in Apps. I and J, we note that the logical basis operators $\sigma^{\nu}_{f_{n}}(t)$ (K.157) still realize the Pauli algebra on site $f_{n}\gg i_{n}$. As before, the strategy is for the pair observables $A_{s,n,x}$ and $A_{s,n,z}$ to be measured on nearby sites $j_{s,n}$ and $j_{s,n}+v$, and then converted into a pair of operators of the same type ($\nu$) under the next unitary channel encountered in the Heisenberg picture—i.e.,
$\bar{A}_{s,n,\nu}\to\sigma^{\nu}_{j_{s,n}}\,\sigma^{\nu}_{j_{s,n}+v}~{}~{}~{}{\rm and}~{}~{}~{}\sigma^{\nu}_{f_{n}}\to\sigma^{\nu}_{f_{n}-v}\,,~{}~{}$ (K.158)
which can always be achieved in duration $\Delta T\geq 1$ Friedman et al. (2022a).
As before, the next step is to send the pair of $\sigma^{\nu}$ operators created on nearby sites via (K.158) in opposite directions. Note that we need to “save” $\Delta T\geq 1$ of the total unitary evolution to “undo” Eq. K.158. The next segment of the Heisenberg evolution of the logical operator (K.157), with duration $T-2$, sends $f_{n}-v\to f_{n}-v(T-1)$, $j_{s,n}\to j_{s,n}-v(T-1)$, and $j_{s,n}+v\to j_{s,n}+v(T-1)$. At this point, the sites of anticommutation must be “undone” (since the initial state is a product state, by Def. 6), which is possible using the inverse of the channel that achieves Eq. K.158. For this to succeed—and for $\sigma^{\nu}_{f_{n}}(T)$ to realize the Pauli algebra on site $i_{n}$—we must have
$d(f_{n},j_{m,n})$ $\leq 2vT-v$ (K.159a)
$d(j_{s+1,n},j_{s,n})$ $\leq 2v(T-1)$ (K.159b)
$d(j_{1,n},i_{n})$ $\leq v(T-1)\,,~{}~{}$ (K.159c)
@@
Finally, we prove Eq. 48. As always, the maximum distance between equivalent sites $j_{s,n}$ and $j_{s+1,n}$ is $\ell\leq 2v(T-1)$. However, we note that at least $2k$ sites must be measured in each region; hence, we must have $\ell\geq 2k$. Combining these inequalities, we find that $2v(T-1)\geq 2k$, which simplifies to Eq. 48. In the case $k=1$, we find $T\geq 2$, as noted in Lemmas 11 and 12. ∎
@@
$\left[A_{j},\,\Gamma(t_{j})\right]=0\,,~{}~{}$ (39)
@@
where $\Gamma(t_{j})$ is some final-state logical operator $\Gamma$ evolved to the (Heisenberg) time $t_{j}$ immediately prior to the measurement of $A_{j}$. The other important result is
@@
$\Gamma(t_{j-1})=\operatorname*{tr}_{{\rm ss}~{}j}\left[\,\mathcal{M}^{\dagger}_{j}\,\Gamma(t_{j})\,\mathcal{M}_{j}\,\right]=\bar{A}^{\lambda_{j,\Gamma}}_{j}\,\Gamma(t_{j})\,,~{}$ (40)
@@
where $\lambda_{j,\Gamma}\in\{0,1\}$ is determined by the recovery channels conditioned on the state of the $j$th Stinespring qubit, and $\bar{A}_{j}$ is the involutory part (4) of the observable $A_{j}$.
The final measured observable $A_{M}$ therefore satisfies
$\left[A_{M},\,\sigma^{\nu}_{f}\right]=0\,,~{}~{}$ (L.161)
@@
for all final logical sites $f\in F$ and all $\nu\in\{x,y,z\}$, which requires that $A_{M}$ has no support in $F$ (since only $\mathds{1}_{f}$ commutes with $\sigma^{\nu}_{f}$ for all nontrivial values of $\nu$).
@@
The penultimate observable $A_{M-1}$ also has no support on any logical site $f\in F$ to which $A_{M}$ is not attached; for any logical operator $\Gamma$ to which $A_{M}$ is attached,
00 0 $=\left[A_{M-1},\,\bar{A}_{M}\,\Gamma\right]$
$=\left[A_{M-1},\,\bar{A}_{M}\right]\,\Gamma+\bar{A}_{M}\,\left[A_{M-1},\,\Gamma\right]\,,~{}$ (L.162)
@@
where at least one of the two commutators in the sum above must vanish, as $A_{M-1}$ (32) remains a single-qubit operator in the canonical-form protocol $\mathcal{W}$ (31) by the requirement of Def. 6 that no unitaries be applied to previously measured sites. Hence, both commutators must vanish, meaning that
$\left[A_{M-1},\,A_{M}\right]=0\,,~{}~{}$ (L.163)
and also that $A_{M-1}$ has no support in $F$.
@@
Repeating this analysis for all subsequent observables $A_{j}$, we again find constraints of the same form as Eq. L.162, with $\bar{A}_{M}$ replaced by some product of observables $\bar{A}_{j+1}$ through $\bar{A}_{M}$. For each $j<M$, we find that none of the measured observables previously encountered in the Heisenberg evolution of $\Gamma$ have support in $F$. As a result, Eq. 39 can be split into a vanishing sum of two commutators à la Eq. L.162, where at least one of the commutators must be zero. Hence, both commutators must vanish, meaning that $A_{j}$ has no support in $F$ and
$\left[A_{j},\,\bar{A}^{\lambda_{j+1,\Gamma}}_{j+1}\cdots\bar{A}^{\lambda_{M,\Gamma}}_{M}\right]=0\,,~{}~{}$ (L.164)
@@
meaning that $A_{j}$ commutes with every product of observables $A_{j+1}$ through $A_{M}$ attached to any logical operator $\Gamma=\sigma^{\nu}_{f}$ via Eq. 40 of Lemma 8.
@@
Importantly, all of these observables are single-qubit operators, and trivially commute unless they act on the same qubit. By Eq. L.164, either all observables correspond to different sites $j\notin F$ or, following the measurement of $A_{j}$ (in the Schrödinger picture), two observables that anticommute with $A_{j}$ are subsequently measured and both are attached to the same logical operators. However, these observables must commute with each other, and hence must be the same, so the product of their involutory parts (4) is the identity. Hence, this pair of measurements can be omitted entirely. Any measurements intermediate to this pair act on other sites, and hence, trivially commute.
@@
This also agrees with the basic intuition that the subsequent anticommuting measurements necessarily result in random outcomes, yet the final logical state must be the same along every outcome trajectory, meaning that no nontrivial operations can be conditioned on the outcomes of the measurements that anticommute with $A_{j}$. More important, however, is the fact that this anticommuting operator is effectively removed from all logical operators prior to encountering $A_{j}$ (in the Heisenberg picture), and hence, these measurements may be omitted entirely without affecting any commutation condition (L.164) on the observables or the evolution of any logical operator. ∎
@@
$\lvert\Psi_{t}\rangle=\mathcal{U}\,\lvert\Psi_{0}\rangle\,,~{}~{}$ (36)
where $\lvert\Psi_{0}\rangle=\lvert\psi\rangle\otimes\lvert\Phi_{0}\rangle$ (6) is the initial state, with $\lvert\Phi_{0}\rangle$ a product state by Def. 3. Importantly, the teleportation conditions (1) of Prop. 1 guarantee that
$\sigma^{\nu}_{i}\,\mathsf{S}_{\nu}=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{W}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{W}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]\,,~{}~{}$ (1)
where $\mathsf{S}_{\nu}\,\lvert\Psi_{0}\rangle=\lvert\Psi_{0}\rangle$ is a stabilizer $\mathsf{S}_{\nu}\in\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$.
@@
$=\bar{A}_{1}^{\lambda_{1,\nu}}\cdots\bar{A}_{M}^{\lambda_{M,\nu}}\,\sigma^{\nu}_{f}=\bar{A}_{\nu}\,\sigma^{\nu}_{f}$
$=\left(\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\right)\,\left(\mathcal{U}\,\mathsf{S}_{\nu}\,\mathcal{U}^{\dagger}\right)\,,~{}~{}$ (M.165)
@@
$\mathsf{S}_{\nu,t}$ $\equiv\mathcal{U}\,\mathsf{S}_{\nu}\,\mathcal{U}^{\dagger}$ (M.166)
which is an element of the (unitary) stabilizer group $\operatorname*{stab}\left(\lvert\Psi_{t}\rangle\right)$ for the resource state, since
$\mathsf{S}_{\nu,t}\,\lvert\Psi_{t}\rangle$ $=\left(\mathcal{U}\,\mathsf{S}_{\nu}\,\mathcal{U}^{\dagger}\right)\,\mathcal{U}\,\lvert\Psi_{0}\rangle$
$=\mathcal{U}\,\mathsf{S}_{\nu}\,\lvert\Psi_{0}\rangle=\mathcal{U}\,\lvert\Psi_{0}\rangle=\lvert\Psi_{t}\rangle\,,~{}~{}$ (M.167)
@@
$\mathsf{S}_{\nu,t}$ $=\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}_{\nu}\,\sigma^{\nu}_{f}\,,~{}~{}$ (M.168)
@@
11 11 $=\left\langle\Psi_{t}\middle|\,\mathsf{S}_{\nu,t}\,\middle|\Psi_{t}\right\rangle$
$=\left\langle\Psi_{t}\middle|\,\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}_{\nu}\,\sigma^{\nu}_{f}\,\middle|\Psi_{t}\right\rangle\,,~{}~{}$ (15)
as stated in the Lemma. The commutation of the “bulk” operator content of the string order parameter $\mathsf{S}_{\nu,t}$ (M.168) is guaranteed by Cor. 14 for all necessary measurements, where we simply omit unnecessary measurements. The anticommutation of the endpoint operators follows automatically from the Pauli algebra. ∎
@@
We first show that there exist subregions of the interval $[i,f]$ with different left endpoints such that a string order parameter analogous to Eq. 15 has unit expectation value in the same resource state $\lvert\Psi_{t}\rangle$ (36). This requires that multiple observables $\bar{A}_{j}$ are attached to both logical operators, which is guaranteed when $L\gg vT$. We show the same for different right endpoints subsequently.
@@
$\sigma^{\nu}_{i}\,\mathsf{S}_{\nu}$ $=\left(\mathcal{U}^{\dagger}\,\bar{A}_{\nu}\,\mathcal{U}\right)\left(\mathcal{U}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{U}\right)\,,~{}~{}$ (N.169)
and importantly, the fact that $L>vT$ (as required for physical teleportation via Def. 3) guarantees that $\mathcal{U}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{U}$ only contributes to the stabilizer part $\mathsf{S}_{\nu}$ of Eq. N.169. Hence, the logical part $\sigma^{\nu}_{i}$ must recover from evolving the leftmost measured observables in $\bar{A}_{\nu}$ (55) under $\mathcal{U}$.
@@
$\bar{A}_{\nu}=\bar{A}_{\nu,1}\cdots\bar{A}_{\nu,m}\,,~{}~{}$ (N.170)
where each operator $\bar{A}_{\nu,s}$ corresponds to a particular “measurement region,” and may represent a product of the involutory parts (4) of two or more measured observables attached to $\sigma^{\nu}_{f}$ via Eq. 40 of Lemma 8.
@@
$\sigma^{\nu}_{i}\,\mathsf{S}_{\nu}$ $=\left(\mathcal{U}^{\dagger}\,\bar{A}_{\nu,1}\,\mathcal{U}\right)\left(\mathcal{U}^{\dagger}\,\bar{A}_{\nu,2}\cdots\bar{A}_{\nu,m}\,\sigma^{\nu}_{f}\,\mathcal{U}\right)\,,~{}$ (N.171)
@@
$\sigma^{\nu}_{i}\,\Sigma^{\nu}_{1}=\mathcal{U}^{\dagger}\,\bar{A}_{\nu,1}\,\mathcal{U}\,,~{}~{}$ (N.172)
where $\Sigma^{\nu}_{1}$ acts nontrivially only to the right of site $i$, when the above is applied to the initial state $\lvert\Psi_{0}\rangle$ (6). Since the right-hand sides of Eq. N.172 for $\nu=x,z$ commute with one another, the left-hand sides must as well. This implies that, e.g., $\Sigma^{x}_{1}$ and $\Sigma^{z}_{1}$ reproduce the Pauli algebra to the right of $i$, up stabilizer elements of $\operatorname*{stab}\left(\lvert\Psi_{0}\rangle\right)$.
The string order parameter (15) from Lemma 15 can be written as the following element of $\operatorname*{stab}\left(\lvert\Psi_{t}\rangle\right)$,
$\mathsf{S}_{\nu,t}=\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}_{\nu}\,\sigma^{\nu}_{f}\,,~{}~{}$ (M.168)
@@
$\mathsf{S}_{\nu,t}$ $=\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}_{\nu,1}\,\left(\mathcal{U}\,\mathcal{U}^{\dagger}\right)\,\bar{A}_{\nu,2}\cdots\bar{A}_{\nu,m}\,\sigma^{\nu}_{f}$
$=\mathcal{U}\,\sigma^{\nu}_{i}\,\left(\mathcal{U}^{\dagger}\,\bar{A}_{\nu,1}\,\mathcal{U}\right)\mathcal{U}^{\dagger}\,\bar{A}_{\nu,2}\cdots\bar{A}_{\nu,m}\,\sigma^{\nu}_{f}$
$=\mathcal{U}\,\sigma^{\nu}_{i}\,\left(\sigma^{\nu}_{i}\,\Sigma^{\nu}_{1}\right)\mathcal{U}^{\dagger}\,\bar{A}_{\nu,2}\cdots\bar{A}_{\nu,m}\,\sigma^{\nu}_{f}$
$=\left(\mathcal{U}\,\Sigma^{\nu}_{1}\,\mathcal{U}^{\dagger}\right)\,\bar{A}_{\nu,2}\cdots\bar{A}_{\nu,m}\,\sigma^{\nu}_{f}\,,~{}~{}$ (N.173)

### v3

@@ Appendix I Proof of Lemma 11
$=\bar{A}^{\vphantom{\prime}}_{1,n,\nu}\cdots\bar{A}^{\vphantom{\prime}}_{m,n,\nu}\,\sigma^{\nu}_{f_{n}}\,,~{}~{}$ (K.157)
where the first label on $\bar{A}^{\,}_{s,n,\nu}$ corresponds to the “measurement region” $s\in[1,m]$ [42], $n$ labels the logical qubit, and $\nu$ labels the logical basis operator (e.g., $\nu\in\{x,z\}$).
We now work out the optimal spacing between measurement sites $j^{\,}_{s,n,\nu}$ for different $s$. Following the proofs in Apps. I and J, we note that the logical basis operators $\sigma^{\nu}_{f_{n}}(t)$ (K.157) still realize the Pauli algebra on site $f_{n}\gg i_{n}$. As before, the strategy is for the pair observables $A^{\,}_{s,n,x}$ and $A^{\,}_{s,n,z}$ to be measured on nearby sites $j^{\,}_{s,n}$ and $j^{\,}_{s,n}+v$, and then converted into a pair of operators of the same type ($\nu$) under the next unitary channel encountered in the Heisenberg picture—i.e.,
$\bar{A}^{\vphantom{\prime}}_{s,n,\nu}\to\sigma^{\nu}_{j^{\vphantom{\prime}}_{s,n}}\,\sigma^{\nu}_{j^{\vphantom{\prime}}_{s,n}+v}~{}~{}~{}{\rm and}~{}~{}~{}\sigma^{\nu}_{f_{n}}\to\sigma^{\nu}_{f_{n}-v}\,,~{}~{}$ (K.158)
which can always be achieved in duration $\Delta T\geq 1$ [42].
As before, the next step is to send the pair of $\sigma^{\nu}$ operators created on nearby sites via (K.158) in opposite directions. Note that we need to “save” $\Delta T\geq 1$ of the total unitary evolution to “undo” Eq. K.158. The next segment of the Heisenberg evolution of the logical operator (K.157), with duration $T-2$, sends $f_{n}-v\to f_{n}-v(T-1)$, $j^{\,}_{s,n}\to j^{\,}_{s,n}-v(T-1)$, and $j^{\,}_{s,n}+v\to j^{\,}_{s,n}+v(T-1)$. At this point, the sites of anticommutation must be “undone” (since the initial state is a product state, by Def. 6), which is possible using the inverse of the channel that achieves Eq. K.158. For this to succeed—and for $\sigma^{\nu}_{f_{n}}(T)$ to realize the Pauli algebra on site $i_{n}$—we must have
$d(f^{\vphantom{\prime}}_{n},j^{\vphantom{\prime}}_{m,n})$ $\leq 2vT-v$ (K.159a)
$d(j^{\vphantom{\prime}}_{s+1,n},j^{\vphantom{\prime}}_{s,n})$ $\leq 2v(T-1)$ (K.159b)
$d(j^{\vphantom{\prime}}_{1,n},i^{\vphantom{\prime}}_{n})$ $\leq v(T-1)\,,~{}~{}$ (K.159c)
@@
Finally, we prove Eq. 48. As always, the maximum distance between equivalent sites $j^{\,}_{s,n}$ and $j^{\,}_{s+1,n}$ is $\ell\leq 2v(T-1)$. However, we note that at least $2k$ sites must be measured in each region; hence, we must have $\ell\geq 2k$. Combining these inequalities, we find that $2v(T-1)\geq 2k$, which simplifies to Eq. 48. In the case $k=1$, we find $T\geq 2$, as noted in Lemmas 11 and 12. ∎
@@
$\left[A^{\vphantom{\prime}}_{j},\,\Gamma(t^{\vphantom{\prime}}_{j})\right]=0\,,~{}~{}$ (39)
@@
where $\Gamma(t^{\,}_{j})$ is some final-state logical operator $\Gamma$ evolved to the (Heisenberg) time $t^{\,}_{j}$ immediately prior to the measurement of $A^{\,}_{j}$. The other important result is
@@
$\Gamma(t^{\vphantom{\prime}}_{j-1})=\operatorname*{tr}_{{\rm ss}~{}j}\left[\,\mathcal{M}^{\dagger}_{j}\,\Gamma(t^{\vphantom{\prime}}_{j})\,\mathcal{M}^{\vphantom{\dagger}}_{j}\,\right]=\bar{A}^{\lambda^{\,}_{j,\Gamma}}_{j}\,\Gamma(t^{\vphantom{\prime}}_{j})\,,~{}$ (40)
@@
where $\lambda^{\,}_{j,\Gamma}\in\{0,1\}$ is determined by the recovery channels conditioned on the state of the $j$th Stinespring qubit, and $\bar{A}^{\,}_{j}$ is the involutory part (33) of the observable $A^{\,}_{j}$.
The final measured observable $A^{\,}_{M}$ therefore satisfies
$\left[A^{\vphantom{\prime}}_{M},\,\sigma^{\nu}_{f}\right]=0\,,~{}~{}$ (L.161)
@@
for all final logical sites $f\in F$ and all $\nu\in\{x,y,z\}$, which requires that $A^{\,}_{M}$ has no support in $F$ (since only $\mathds{1}_{f}$ commutes with $\sigma^{\nu}_{f}$ for all nontrivial values of $\nu$).
@@
The penultimate observable $A^{\,}_{M-1}$ also has no support on any logical site $f\in F$ to which $A^{\,}_{M}$ is not attached; for any logical operator $\Gamma$ to which $A^{\,}_{M}$ is attached,
00 0 $=\left[A^{\vphantom{\prime}}_{M-1},\,\bar{A}^{\vphantom{\prime}}_{M}\,\Gamma\right]$
$=\left[A^{\vphantom{\prime}}_{M-1},\,\bar{A}^{\vphantom{\prime}}_{M}\right]\,\Gamma+\bar{A}^{\vphantom{\prime}}_{M}\,\left[A^{\vphantom{\prime}}_{M-1},\,\Gamma\right]\,,~{}$ (L.162)
@@
where at least one of the two commutators in the sum above must vanish, as $A^{\,}_{M-1}$ (32) remains a single-qubit operator in the canonical-form protocol $\mathcal{W}$ (31) by the requirement of Def. 6 that no unitaries be applied to previously measured sites. Hence, both commutators must vanish, meaning that
$\left[A^{\,}_{M-1},\,A^{\,}_{M}\right]=0\,,~{}~{}$ (L.163)
and also that $A^{\,}_{M-1}$ has no support in $F$.
@@
Repeating this analysis for all subsequent observables $A^{\,}_{j}$, we again find constraints of the same form as Eq. L.162, with $\bar{A}^{\,}_{M}$ replaced by some product of observables $\bar{A}^{\vphantom{\prime}}_{j+1}$ through $\bar{A}^{\vphantom{\prime}}_{M}$. For each $j<M$, we find that none of the measured observables previously encountered in the Heisenberg evolution of $\Gamma$ have support in $F$. As a result, Eq. 39 can be split into a vanishing sum of two commutators à la Eq. L.162, where at least one of the commutators must be zero. Hence, both commutators must vanish, meaning that $A^{\,}_{j}$ has no support in $F$ and
$\left[A^{\vphantom{\prime}}_{j},\,\bar{A}^{\lambda_{j+1,\Gamma}}_{j+1}\cdots\bar{A}^{\lambda_{M,\Gamma}}_{M}\right]=0\,,~{}~{}$ (L.164)
@@
meaning that $A^{\,}_{j}$ commutes with every product of observables $A^{\,}_{j+1}$ through $A^{\,}_{M}$ attached to any logical operator $\Gamma=\sigma^{\nu}_{f}$ via Eq. 40 of Lemma 8.
@@
Importantly, all of these observables are single-qubit operators, and trivially commute unless they act on the same qubit. By Eq. L.164, either all observables correspond to different sites $j\notin F$ or, following the measurement of $A^{\,}_{j}$ (in the Schrödinger picture), two observables that anticommute with $A^{\,}_{j}$ are subsequently measured and both are attached to the same logical operators. However, these observables must commute with each other, and hence must be the same, so the product of their involutory parts (33) is the identity. Hence, this pair of measurements can be omitted entirely. Any measurements intermediate to this pair act on other sites, and hence, trivially commute.
@@
This also agrees with the basic intuition that the subsequent anticommuting measurements necessarily result in random outcomes, yet the final logical state must be the same along every outcome trajectory, meaning that no nontrivial operations can be conditioned on the outcomes of the measurements that anticommute with $A^{\,}_{j}$. More important, however, is the fact that this anticommuting operator is effectively removed from all logical operators prior to encountering $A^{\,}_{j}$ (in the Heisenberg picture), and hence, these measurements may be omitted entirely without affecting any commutation condition (L.164) on the observables or the evolution of any logical operator. ∎
@@
$\lvert\Psi^{\vphantom{\prime}}_{t}\rangle=\mathcal{U}\,\lvert\Psi^{\vphantom{\prime}}_{0}\rangle\,,~{}~{}$ (36)
where $\lvert\Psi^{\,}_{0}\rangle=\lvert\psi\rangle\otimes\lvert\Phi^{\,}_{0}\rangle$ (6) is the initial state, with $\lvert\Phi^{\,}_{0}\rangle$ a product state by Def. 3. Importantly, the teleportation conditions (8) of Prop. 1 guarantee that
$\sigma^{\nu}_{i}\,\mathsf{S}^{\vphantom{\prime}}_{\nu}=\operatorname*{tr}_{\rm ss}\left[\,\mathcal{W}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{W}\,\widetilde{P}^{(\boldsymbol{0})}_{\rm ss}\,\right]\,,~{}~{}$ (8)
where $\mathsf{S}^{\,}_{\nu}\,\lvert\Psi^{\,}_{0}\rangle=\lvert\Psi^{\,}_{0}\rangle$ is a stabilizer $\mathsf{S}^{\,}_{\nu}\in\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)$.
@@
$=\bar{A}_{1}^{\lambda^{\,}_{1,\nu}}\cdots\bar{A}_{M}^{\lambda^{\,}_{M,\nu}}\,\sigma^{\nu}_{f}=\bar{A}^{\vphantom{\prime}}_{\nu}\,\sigma^{\nu}_{f}$
$=\left(\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\right)\,\left(\mathcal{U}\,\mathsf{S}^{\vphantom{\prime}}_{\nu}\,\mathcal{U}^{\dagger}\right)\,,~{}~{}$ (M.165)
@@
$\mathsf{S}^{\vphantom{\prime}}_{\nu,t}$ $\equiv\mathcal{U}\,\mathsf{S}^{\,}_{\nu}\,\mathcal{U}^{\dagger}$ (M.166)
which is an element of the (unitary) stabilizer group $\operatorname*{stab}\left(\lvert\Psi^{\,}_{t}\rangle\right)$ for the resource state, since
$\mathsf{S}^{\vphantom{\prime}}_{\nu,t}\,\lvert\Psi^{\vphantom{\prime}}_{t}\rangle$ $=\left(\mathcal{U}\,\mathsf{S}^{\vphantom{\prime}}_{\nu}\,\mathcal{U}^{\dagger}\right)\,\mathcal{U}\,\lvert\Psi^{\vphantom{\prime}}_{0}\rangle$
$=\mathcal{U}\,\mathsf{S}^{\vphantom{\prime}}_{\nu}\,\lvert\Psi^{\vphantom{\prime}}_{0}\rangle=\mathcal{U}\,\lvert\Psi^{\vphantom{\prime}}_{0}\rangle=\lvert\Psi^{\vphantom{\prime}}_{t}\rangle\,,~{}~{}$ (M.167)
@@
$\mathsf{S}^{\vphantom{\prime}}_{\nu,t}$ $=\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu}\,\sigma^{\nu}_{f}\,,~{}~{}$ (M.168)
@@
11 11 $=\left\langle\Psi^{\vphantom{\prime}}_{t}\middle|\,\mathsf{S}^{\vphantom{\prime}}_{\nu,t}\,\middle|\Psi^{\vphantom{\prime}}_{t}\right\rangle$
$=\left\langle\Psi^{\vphantom{\prime}}_{t}\middle|\,\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu}\,\sigma^{\nu}_{f}\,\middle|\Psi^{\vphantom{\prime}}_{t}\right\rangle\,,~{}~{}$ (54)
as stated in the Lemma. The commutation of the “bulk” operator content of the string order parameter $\mathsf{S}^{\,}_{\nu,t}$ (M.168) is guaranteed by Cor. 14 for all necessary measurements, where we simply omit unnecessary measurements. The anticommutation of the endpoint operators follows automatically from the Pauli algebra. ∎
@@
We first show that there exist subregions of the interval $[i,f]$ with different left endpoints such that a string order parameter analogous to Eq. 54 has unit expectation value in the same resource state $\lvert\Psi^{\,}_{t}\rangle$ (36). This requires that multiple observables $\bar{A}^{\,}_{j}$ are attached to both logical operators, which is guaranteed when $L\gg vT$. We show the same for different right endpoints subsequently.
@@
$\sigma^{\nu}_{i}\,\mathsf{S}^{\vphantom{\prime}}_{\nu}$ $=\left(\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu}\,\mathcal{U}\right)\left(\mathcal{U}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{U}\right)\,,~{}~{}$ (N.169)
and importantly, the fact that $L>vT$ (as required for physical teleportation via Def. 3) guarantees that $\mathcal{U}^{\dagger}\,\sigma^{\nu}_{f}\,\mathcal{U}$ only contributes to the stabilizer part $\mathsf{S}^{\,}_{\nu}$ of Eq. N.169. Hence, the logical part $\sigma^{\nu}_{i}$ must recover from evolving the leftmost measured observables in $\bar{A}^{\,}_{\nu}$ (55) under $\mathcal{U}$.
@@
$\bar{A}^{\vphantom{\prime}}_{\nu}=\bar{A}^{\vphantom{\prime}}_{\nu,1}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,m}\,,~{}~{}$ (N.170)
where each operator $\bar{A}^{\,}_{\nu,s}$ corresponds to a particular “measurement region,” and may represent a product of the involutory parts (33) of two or more measured observables attached to $\sigma^{\nu}_{f}$ via Eq. 40 of Lemma 8.
@@
$\sigma^{\nu}_{i}\,\mathsf{S}^{\vphantom{\prime}}_{\nu}$ $=\left(\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,1}\,\mathcal{U}\right)\left(\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,2}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,m}\,\sigma^{\nu}_{f}\,\mathcal{U}\right)\,,~{}$ (N.171)
@@
$\sigma^{\nu}_{i}\,\Sigma^{\nu}_{1}=\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,1}\,\mathcal{U}\,,~{}~{}$ (N.172)
where $\Sigma^{\nu}_{1}$ acts nontrivially only to the right of site $i$, when the above is applied to the initial state $\lvert\Psi^{\,}_{0}\rangle$ (6). Since the right-hand sides of Eq. N.172 for $\nu=x,z$ commute with one another, the left-hand sides must as well. This implies that, e.g., $\Sigma^{x}_{1}$ and $\Sigma^{z}_{1}$ reproduce the Pauli algebra to the right of $i$, up stabilizer elements of $\operatorname*{stab}\left(\lvert\Psi^{\,}_{0}\rangle\right)$.
The string order parameter (54) from Lemma 15 can be written as the following element of $\operatorname*{stab}\left(\lvert\Psi^{\,}_{t}\rangle\right)$,
$\mathsf{S}^{\vphantom{\prime}}_{\nu,t}=\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu}\,\sigma^{\nu}_{f}\,,~{}~{}$ (M.168)
@@
$\mathsf{S}^{\vphantom{\prime}}_{\nu,t}$ $=\mathcal{U}\,\sigma^{\nu}_{i}\,\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,1}\,\left(\mathcal{U}\,\mathcal{U}^{\dagger}\right)\,\bar{A}^{\vphantom{\prime}}_{\nu,2}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,m}\,\sigma^{\nu}_{f}$
$=\mathcal{U}\,\sigma^{\nu}_{i}\,\left(\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,1}\,\mathcal{U}\right)\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,2}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,m}\,\sigma^{\nu}_{f}$
$=\mathcal{U}\,\sigma^{\nu}_{i}\,\left(\sigma^{\nu}_{i}\,\Sigma^{\nu}_{1}\right)\mathcal{U}^{\dagger}\,\bar{A}^{\vphantom{\prime}}_{\nu,2}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,m}\,\sigma^{\nu}_{f}$
$=\left(\mathcal{U}\,\Sigma^{\nu}_{1}\,\mathcal{U}^{\dagger}\right)\,\bar{A}^{\vphantom{\prime}}_{\nu,2}\cdots\bar{A}^{\vphantom{\prime}}_{\nu,m}\,\sigma^{\nu}_{f}\,,~{}~{}$ (N.173)